The larger of two numbers is nine more than four times the smaller number. The sum of the two numbers is sixty-nine. Find the two numbers.
Answer by ikleyn(52780)

.
The larger of two numbers is nine more than four times the smaller number.
The sum of the two numbers is sixty-nine. Find the two numbers.
~~~~~~~~~~~~~~

Let x be the smaller number.

Then the larger number is 4x+9, according to the condition.


The sum of the numbers is 69


    x + (4x+9) = 69.


Simplify and find x


    5x + 9 = 69

    5x     = 69 - 9 = 60

     x              = 60/5 = 12.


ANSWER.  The smaller number is 12.  The greater number is  4*12 + 9 = 48+9 = 57.


CHECK.   12 + 57 = 69,   correct (!)

Solved.